OpenStudy (anonymous):

Need a double check:

OpenStudy (anonymous):

\[\int\limits_{1}^{3} \frac{ 1 }{ x }dx:\] My answer is log(3)-log(1)= 0.477121254...

OpenStudy (anonymous):

Is this the correct assessment?

OpenStudy (anonymous):

use naural logarithm. ln(3)-ln(1)

OpenStudy (anonymous):

Ahh thanks, it's now 1.098612289

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Isn't it ln(1)-ln(3)?

OpenStudy (anonymous):

nope. top minus bottom

OpenStudy (anonymous):

I also have this one:

OpenStudy (anonymous):

\[\int\limits_{-3}^{-1} \frac{ 1 }{ x }dx\] Which when I do ln(-3)-ln(-1) gives me Math error :S

OpenStudy (anonymous):

errr ln(-1)-ln(-3) I mean.

OpenStudy (anonymous):

use this identity: lnA-lnB=ln(A/B) thus ln(-1)-ln(-3)=ln(-1/-3)=ln(1/3)=-1.0986

OpenStudy (anonymous):

That makes sense, I forgot about identities, felt like I was missing something. Thanks! :)

OpenStudy (anonymous):

np